OpenStudy (anonymous):

A manufacturer of salad dressings uses machines to dispense liquid ingredients into bottles that move along a filling line. The machine that dispenses dressings is working properly when the mean amount dispensed is 330 mL. The population standard deviation of the amount dispensed is 4 mL. A sample of 50 bottles is selected periodically, and the filling line is stopped if there is evidence that the mean amount dispensed is different from 330 mL.

OpenStudy (mertsj):

Is there a question hidden in there somewhere?

OpenStudy (anonymous):

i beleive so...

OpenStudy (anonymous):

Alternate Hypothesis H1: μ ≠ μ0 (Alternate Hypothesis) is true with at least 95% confidence. Yes; the sample shows a different average value than 330mL

OpenStudy (anonymous):

thanks mate, would it be possible if you could shhow your working? im struggling with how to use the formula. thankyou

OpenStudy (anonymous):

SINGLE SAMPLE TEST, TWO-TAILED, there's 6 - Step Procedure for t Distributions, "two-tailed test"

OpenStudy (anonymous):

Step 1: State the hypothesis to be tested. Null Hypothesis H0: μ = μ0 Alternate Hypothesis H1: μ ≠ μ0 Step 2: Determine a planning value for α [level of significance] = 0.05 Step 3: From the sample data determine x-bar, s and n; then compute Standardized Test Statistic: t = ( x-bar - μ0 )/( s/ SQRT(n) ) x-bar: Est. of the Pop. Mean (statistical mean of the sample) = 329.7 n: number of individuals in the sample = 50 s: sample standard deviation = 0.566 μ0: Population Mean = 330 Standardized Test Statistic t = ( 329.7 - 330 )/( 0.566 / SQRT( 50 )) = 3.748 Step 4: Using Students t distribution, 'lookup' the area outside of t = TDIST( 3.748 , 49 , 2 ) Step 5: Area in Step 4 is equal to P value = 0 based on n -1 = 49 df (degrees of freedom). Table look-up value shows area under the 49 df curve outside of t = +/- 3.748 is (approx.) P value = 0 [2 * 0.0002] by addition of both 'tails' of t distribution. Step 6: For P ≥ α, fail to reject H0; and for P < α, reject H0 with 95% confidence in the conclusion. Conclusion: Alternate Hypothesis H1: μ ≠ μ0 (Alternate Hypothesis) is true with at least 95% confidence. Note: level of significance [α] is the maximum level of risk an experimenter is willing to take in making a "reject H0" or "conclude H1" conclusion (i.e. it is the maximum risk in making a Type I error).

OpenStudy (anonymous):

thankyou very much Shruti, much appreciated

OpenStudy (anonymous):

not much..coz you have'nt given me any medal at all =D